OpenStudy (anonymous):

Solve 6x^2+x-12=0 by factoring. write the trinomial as a four term polynomial and the factor.

jhonyy9 (jhonyy9):

x_1,_2=(-1+/-sqrt 1+288 )/12 = (-1+/- 17)/12 = -18/12 = -3/2 and 16/12 = =4/3 (x+3/2) *( x-4/3) (2x+3)* (3x-4)

jhonyy9 (jhonyy9):

\[6x ^{2}+x-12=(2x+3)(3x-4)\]